# Dataset and Assessment of LLMs for Cultural Text Adaptation in Indian Context

### v1

@@ D.2 Significance Testing
For every dataset and model, we report the mean exact-match accuracy and a 95% confidence interval. We use a non-parametric bootstrap Efron (1979) with 10,000 resamples. Each resample picks $n$ examples with replacement from the test set, we recompute the mean, and we take the 2.5–97.5 percentile range as the 95% CI. When comparing two systems on the same items, we use a paired approximate randomisation test (10,000 sign flips; Dror et al. (2018)). We flip the sign of each per-item difference at random and recompute the mean difference. The two-sided $p$-value is the fraction of permutations with an absolute difference at least as large as the observed one (with the usual +1 correction). We also report the mean gap in percentage points (pp).
@@
Within each family (all model–model pairs on a given dataset/prompt, and all DIWALI–CANDLE deltas per model on that dataset/prompt), we apply Benjamini–Hochberg (BH) Benjamini and Hochberg (2018). We call a result significant if $q<0.05$.

### v2

@@ D.2 Significance Testing
For every dataset and model, we report the mean exact-match accuracy and a 95% confidence interval. We use a non-parametric bootstrap Efron1979BootstrapMA with 10,000 resamples. Each resample picks $n$ examples with replacement from the test set, we recompute the mean, and we take the 2.5–97.5 percentile range as the 95% CI. When comparing two systems on the same items, we use a paired approximate randomisation test (10,000 sign flips; dror-etal-2018-hitchhikers). We flip the sign of each per-item difference at random and recompute the mean difference. The two-sided $p$-value is the fraction of permutations with an absolute difference at least as large as the observed one (with the usual +1 correction). We also report the mean gap in percentage points (pp).
@@
Within each family (all model–model pairs on a given dataset/prompt, and all DIWALI–CANDLE deltas per model on that dataset/prompt), we apply Benjamini–Hochberg (BH) 10.1111/j.2517-6161.1995.tb02031.x. We call a result significant if $q<0.05$.